OpenStudy (anonymous):

Express answer in exact form. A regular hexagon with sides of 3" is inscribed in a circle. Find the area of a segment formed by a side of the hexagon and the circle. (Hint: remember Corollary 1--the area of an equilateral triangle is 1/4 s2 √3.) help me please

OpenStudy (anonymous):

area segment = area sector - area triangle

OpenStudy (anonymous):

area of sector = 1/6 pi r^2 r = 3

OpenStudy (anonymous):

height of triangle = sqrt(3^2 - 1.5^2 ) = 3sqrt3 / 2

OpenStudy (anonymous):

sorry - i have to go